# Online Learning of Whittle Indices for Restless Bandits with Non-Stationary Transition Kernels

### v1

@@ Abstract
We consider optimal resource allocation for restless multi-armed bandits (RMABs) in unknown, non-stationary settings. RMABs are PSPACE-hard to solve optimally, even when all parameters are known. The Whittle index policy is known to achieve asymptotic optimality for a large class of such problems, while remaining computationally efficient. In many practical settings, however, the transition kernels required to compute the Whittle index are unknown and non-stationary. In this work, we propose an online learning algorithm for Whittle indices in this setting. Our algorithm first predicts current transition kernels by solving a linear optimization problem based on upper confidence bounds and empirical transition probabilities calculated from data over a sliding window. Then, it computes the Whittle index associated with the predicted transition kernels. We design these sliding windows and upper confidence bounds to guarantee sub-linear dynamic regret on the number of episodes $T$, under the condition that transition kernels change slowly over time (rate upper bounded by $\epsilon=1/T^{k}$ with $k>0$). Furthermore, our proposed algorithm and regret analysis are designed to exploit prior domain knowledge and structural information of the RMABs to accelerate the learning process. Numerical results validate that our algorithm achieves superior performance in terms of lowest cumulative regret relative to baselines in non-stationary environments.
@@ 1 Introduction
Even when the transition kernels are known, developing an optimal policy for RMABs is PSPACE-hard [18] because of the combinatorial action space. The Whittle index policy [1] achieves asymptotic optimality [19, 20] for a special class of RMABs that satisfy the indexability property. Most prior works in in this space focus on stationary transition kernels (either known [21, 4, 5, 22, 12] or unknown [23, 24]), leaving a large gap for studying RMABs with unknown and non-stationary transition kernels. In addition, it is common in many applications to have prior knowledge regarding the transition kernels for some parts of the state space. For example, consider a wireless scheduling problem which aims to maximize information freshness in selecting which users (arms) to schedule. In this case, the state can be modeled using Age of Information (AoI) [25, 26] – a widely used metric for quantifying information freshness. Then, the AoI of an arm increases by one if the arm is not scheduled for transmission. Conversely, if the arm is scheduled, its AoI resets to one with the success probability of the transmission. Thus, the AoI will never increase by 2 or decrease to a value other than 1. Applying techniques from online and reinforcement learning naively would require attempting to learn the entire transition matrix, even though parts of it are known, leading to poor performance and inefficient regret guarantees.
In this paper, we pose the following research question: Can we develop a Whittle index-based online algorithm for RMABs with non-stationary transition kernels? To address this, we take prior domain knowledge for RMABs into account and provide online algorithms and regret guarantees that scale with the amount of uncertainty about the bandits.
@@
  • We model non-stationarity of RMABs by assuming an arbitrary upper-bound $\epsilon$ on the total variational divergence between transition kernels in two consecutive episodes. Based on this, we propose an online learning algorithm which predicts the Whittle index by using a sliding window and upper confidence bound approaches. The sliding window and the upper confidence bound are determined as functions of the parameter $\epsilon$. Moreover, our solution incorporates prior domain knowledge and structural information of the RMABs. In particular, our algorithm employs different window sizes and upper confidence bounds for the subset of state-action pairs with unknown yet stationary transition kernels compared to state-action pairs with non-stationary transition kernels.
  • We rigorously characterize an upper bound on the dynamic regret of our algorithm. Our analysis demonstrates the impact of the parameter $\epsilon$ and the window size $w$ on the dynamic regret, allowing design of $w$ as a function of $\epsilon$. Given $\epsilon\leq 1/T^{k}$, our design ensures sub-linear dynamic regret $O(\sqrt{\mathrm{log}(T)}T^{r})$ on the total number of episodes $T$ with $k>0$, $r=\max\{(3-k)/3,1/2\}$, and $w=T^{q}$ with $q=\min\{2k/3,1\}$. In addition, our results provide tighter regret bounds for many applications where system designers possess some knowledge of the transition kernel structure beforehand and can be effectively to problems exhibiting stationary and non-stationary dynamics.
  • Our simulation results show that our algorithm achieves sub-linear regret in the total number of episodes. They also show that our algorithm achieves lower regret in practice compared with the UCWhittle policy [23], WIQL policy [27], and a uniformly randomized policy [3] baselines.
@@ 2 Related Work
Offline Whittle Index Policy for RMABs: Whittle’s seminal work [1] introduced a heuristic policy for the infinite-horizon RMAB problem, known as the Whittle index policy. This policy relies on establishing a special mathematical property called indexability for each arm and then deriving functions called index functions that map states to how valuable it would be to activate an arm at that state. Running the policy simply requires activating the $M$ bandits with the highest Whittle indices out of the $N$ bandits at each decision time. The Whittle index achieves asymptotic optimality [19, 20] when the indexability condition [1, 28] is satisfied. Motivated by Whittle’s work, many subsequent works have applied the Whittle index framework to different resource allocation problems [21, 4, 5, 22, 12, 3, 6, 29] by modeling them as RMABs.
Online Learning of Whittle Index: When the transition kernels of a RMAB are unknown and stationary, the problem of finding the Whittle index becomes an online/reinforcement learning problem. In this effort, multiple works [30, 31, 27] have proposed Q-learning algorithms to compute Whittle Index. Authors in [32] proposed NeurWIN and [33] proposed DeepTOP to compute Whittle index by using neural networks. These prior works did not provide any regret guarantees for their policy. Evaluating regret with respect to the optimal policy requires computing the optimal solution to the RMAB problem, which is intractable. In [34] the authors develop an online Whittle algorithm with static regret guarantees compared to the best fixed Whittle index policy. However, they only consider time-varying rewards, the transition kernels remain static. Another approach to make the objective tractable is proposed in [23], which defines regret in terms of the relaxed Lagrangian of the RMAB. In this Lagrangian setting, the authors of [23] present UCWhittle, a Whittle index-based algorithm that uses an upper confidence bound approach to learn the Whittle index. The UCWhittle algorithm achieves $O(\sqrt{T\mathrm{log}T})$ for RMAB with unknown and stationary transition kernels. Unlike prior works, we propose a Whittle index-based algorithm that uses a sliding window and an upper confidence bound approach to learn Whittle index for non-stationary transition dynamics, with provable regret bounds. To the best of our knowledge, this is the first work to provide dynamic regret analysis of an online Whittle index-based policy for RMABs with unknown and non-stationary transition dynamics.
@@ 3 Problem Setting
We consider an episodic RMAB problem with $N$ arms and an unknown non-stationary environment. Each arm $n\in[N]$ is associated with a unichain MDP denoted by a tuple $(\mathcal{S},\mathcal{A},P_{n,t},r_{n})$ at every episode $t$, where the state space $\mathcal{S}$ is finite, $\mathcal{A}=\{0,1\}$ is a set of binary actions, $P_{n,t}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ is the transition kernel of arm $n$ with $P_{n,t}(s^{\prime}_{n}|s_{n},a_{n})$ being the probability of transitioning to state $s^{\prime}_{n}$ from state $s_{n}$ by taking action $a_{n}$ in episode $t$, and $r_{n}(s_{n},a_{n})$ is the reward function for arm $n$ when the current state is $s_{n}$ and the action $a_{n}$ is taken. The total number of episodes is $T$ and each episode itself consists of $H$ time slots. We consider that $P_{n,t}$ is unknown and non-stationary, i.e., $P_{n,t}$ can change over episode $t\in[T]$.
A decision maker (DM) determines what action to apply to each arm at a decision time $h\in[H]$ of an episode $t\in[T]$ under the instantaneous activation constraint that at most $M$ arms can be activated. The action taken by the DM in episode $t$ is described by a deterministic policy $\pi_{t}:\mathcal{S}^{N}\mapsto\mathcal{A}^{N}$ which maps a given state $\mathbf{s}=(s_{1},s_{2},\ldots,s_{M})\in\mathcal{S}^{N}$ to an action $\mathbf{a}=(a_{1},a_{2},\ldots,a_{M})\in\mathcal{A}^{N}$. The corresponding expected discounted reward in episode $t$ is given by
@@
where $\Pi$ is the set of all causal policy $\pi_{t}:\mathcal{S}^{N}\mapsto\{0,1\}^{N}$ that maps the state $\mathbf{s}\in\mathcal{S}^{N}$ to $\mathbf{a}\in\mathcal{A}^{N}$.
@@ 3.1 Lagrangian Relaxation
Because the main problem (2) is intractable, we relax the per-time slot constraint and use the Lagrangian defined below:
@@
The Lagrangian problem (3) enables us to decompose the combinatorial decision problem (2) into a set of $N$ independent Markov decision process for each arm:
$U\left(\pi^{*}_{n,t},P_{n,t},\lambda\right)=\max_{\pi_{n,t}\in\Pi_{n}}\mathbb{E}\left[\sum_{h=1}^{H}\gamma^{h-1}\bigg{(}r_{n}(s_{n,h,t},a_{n,h,t})-\lambda a_{n,h,t}\bigg{)}\bigg{|}\pi_{n,t},P_{n,t}\right],$ (4)
where we optimize policy $\pi_{n,t}:\mathcal{S}\mapsto\mathcal{A}$ for each arm $n$ from the set of all causal policies $\pi_{n,t}\in\Pi_{n}$.
@@ 3.2 Whittle Index Policy
Given $\lambda$, we denote $\phi_{n}(\lambda)$ by the set of states for which it is optimal to activate the arm. The set $\phi_{n}(\lambda)$ is given by $\phi_{n}(\lambda):=\{s\in\mathcal{S}:Q_{n,\lambda,t}(s,0)>Q_{n,\lambda,t}(s,1)\}$, where the action value function $Q_{n,\lambda,t}(s,a)$ associated with Bellman optimality equation for (4) is
@@
and the value function $V_{n,\lambda,t}(s)$ associated with Bellman optimality equation for (4) is
@@
Definition 1 (Indexability).
An arm is said to be indexable if the set $\phi_{n}(\lambda)$ increases monotonically as $\lambda$ increases from 00 to $\infty$. A restless bandit problem is said to be indexable if all arms are indexable.
Definition 2 (Whittle Index).
@@
Whittle Index Policy activates at most $M$ arms out of $N$ arms with highest Whittle indices. However, as we can observe from (7), we can compute Whittle index if we know the transition kernel $P_{n,t}$ of every episode $t\in[T]$. Next, we model how transition kernels change over every episode.
3.3 Transition Kernels Model
In this section, we model the non-stationarity of transition kernels in RMABs. Since transition kernels in an RMAB may drift at varying rates across different arms $n\in[N]$, we define an arm-specific upper bound, $\epsilon_{n}$, on the rate at which the transition kernel of arm $n$ can change over two consecutive episodes.
We consider that the total variation distance [35] between transition kernels in two consecutive episodes is upper bounded by a parameter $\epsilon_{n}$ for arm $n$. Specifically, we have
$\max_{(s,a)\in\mathcal{S}\times A}\sum_{s^{\prime}\in\mathcal{S}}\|P_{n,t}(s^{\prime}|s,a)-P_{n,t-1}(s^{\prime}|s,a)\|_{1}\leq\epsilon_{n}.$ (8)
The parameter $\epsilon_{n}$ determines how quickly the transition kernels change. If $\epsilon_{n}\rightarrow 0$, then transition kernels of arm $n$ are close to stationary, whereas if $\epsilon_{n}$ is far from 00, the transition kernels of arm $n$ become highly non-stationary.
In an RMAB problem, stationary and non-stationary transition kernels can co-exist. To address this co-existing dynamics, we categorize the state-action pairs for each arm $n\in[N]$ into the following two sets based on prior domain knowledge:
$\mathcal{Z}_{1}^{(n)}$ =$=\{(s,a)\in\mathcal{S}\times\mathcal{A}:P_{n,t}(s^{\prime}|s,a)=q(s^{\prime},a,s)~{}\text{is stationary but unknown},~{}\forall s^{\prime}\in\mathcal{S}\},$ (9)
$\mathcal{Z}_{2}^{(n)}$ =$=\{(s,a)\in\mathcal{S}\times\mathcal{A}:P_{n,t}(s^{\prime}|s,a)~{}\text{is non-stationary and unknown},~{}\forall s^{\prime}\in\mathcal{S}\}.$ (10)
If $|\mathcal{Z}^{(n)}_{1}|=0$ and $|\mathcal{Z}^{(n)}_{2}|=|\mathcal{S}\times\mathcal{A}|$, then the transition probabilities for all state-action pairs in arm $n$ are non-stationary. On the other hand, If $|\mathcal{Z}^{(n)}_{1}|=|\mathcal{S}\times\mathcal{A}|$ and $|\mathcal{Z}^{(n)}_{2}|=0$, then the transition probabilities for all state-action pairs in arm $n$ are stationary.
In many applications, the probability transition kernels are sparse - meaning that many state transitions are not possible under certain actions. To model this we introduce $\mathcal{S}_{0}^{(n)}(s,a)$ as the set of all states $s^{\prime}\in\mathcal{S}$ such that the probability to transit from state $s\in\mathcal{S}$ to state $s^{\prime}\in\mathcal{S}$ given action $a\in\mathcal{A}$ is always 00, i.e.,
$\mathcal{S}_{0}^{(n)}(s,a)$ =$=\{s^{\prime}\in\mathcal{S}:P_{n,t}(s^{\prime}|s,a)=0,\forall t\}.$ (11)
The sets $\mathcal{S}_{0}^{(n)}(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$ represents the sparsity of transition kernels for arm $n$.
3.4 Problem Statement
The DM is assumed to know the parameter $\epsilon_{n}$, the sets $\mathcal{Z}^{(n)}_{1}$, $\mathcal{Z}^{(n)}_{2}$, and $\mathcal{S}_{0}^{(n)}(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. The goal of the DM is to design an online algorithm that runs over a sequence of episodes and maximizes the sum of discounted rewards across all episodes. Towards this end, we develop our online algorithm that learns the probability transition kernels and uses the Whittle Index to pick approximately optimal policies in each episode. In the next section, we provide our online algorithm.

### v2

@@ Abstract
We study optimal resource allocation in restless multi-armed bandits (RMABs) under unknown and non-stationary dynamics. Solving RMABs optimally is PSPACE-hard even with full knowledge of model parameters, and while the Whittle index policy offers asymptotic optimality with low computational cost, it requires access to stationary transition kernels - an unrealistic assumption in many applications. To address this challenge, we propose a Sliding-Window Online Whittle (SW-Whittle) policy that remains computationally efficient while adapting to time-varying kernels. Our algorithm achieves a dynamic regret of $\tilde{O}(T^{2/3}\tilde{V}^{1/3}+T^{4/5})$ for large RMABs, where $T$ is the number of episodes and $\tilde{V}$ is the total variation distance between consecutive transition kernels. Importantly, we handle the challenging case where the variation budget is unknown in advance by combining a Bandit-over-Bandit framework with our sliding-window design. Window lengths are tuned online as a function of the estimated variation, while Whittle indices are computed via an upper-confidence-bound of the estimated transition kernels and a bilinear optimization routine. Numerical experiments demonstrate that our algorithm consistently outperforms baselines, achieving the lowest cumulative regret across a range of non-stationary environments.
@@ 1 Introduction
When the transition kernels of a RMAB are unknown and non-stationary, the problem of finding the Whittle index becomes an online/reinforcement learning problem. Many papers designed algorithms for MDPs and MAB with unknown and non-stationary transition kernels and analyzed dynamic regret for MDPs and MAB [18, 19, 20, 21, 22]. However, algorithm designed in these prior works can not be applied to RMAB due its special structures: The passive arms (arms that are not activated) continue to evolve stochastically. Because of the special structure and the combinatorial action space, even when the transition kernels are known, developing an optimal policy for RMABs is PSPACE-hard [23]. Whittle’s seminal work [1] introduced a heuristic policy for RMAB problem, known as the Whittle index policy. This policy relies on establishing a special mathematical property called indexability for each arm and then deriving functions called index functions that map states to how valuable it would be to activate an arm at that state. Running the policy simply requires activating the $M$ bandits with the highest Whittle indices out of the $N$ bandits at each decision time. To compute Whittle index, the problem is decomposed to multiple single-arm MDPs after a Lagrangian relaxation technique. Then, Whittle index is computed using the solution of the multiple MDPs. The Whittle index achieves asymptotic optimality, if the RMAB is indexable and has a global attractor point [24, 25, 26, 27]. Most prior works in utilizing Whittle index-based policy focus on known and stationary transition kernels [28, 4, 5, 29, 12].
Applying traditional online learning and reinforcement learning policies [18, 19, 20, 21, 22] naively to RMAB with unknown and non-stationary transition kernels may lead to inefficient learning performance and to exponential regret bounds. This necessitates combining Whittle with online learning methods. In this direction, a recent work [30]) designed a Whittle index-based policy called UCWhittle for unknown but stationary transition kernels. Although techniques exist for adapting reinforcement learning algorithms to non-stationary environments [22], they are not directly compatible with the recently proposed UCWhittle policy [30]. This incompatibility arises from the unique structure of RMABs and the specific method used to compute the Whittle index via Lagrangian relaxation.
In addition, it is common in many applications to have prior knowledge regarding the sparsity of transition kernels for some parts of the state space. For example, consider a wireless scheduling problem which aims to maximize information freshness in selecting which users (arms) to schedule. In this case, the state can be modeled using Age of Information (AoI) [31, 32] – a widely used metric for quantifying information freshness. Then, the AoI of an arm increases by one if the arm is not scheduled for transmission. Conversely, if the arm is scheduled, its AoI resets to one with the success probability of the transmission. Thus, the AoI will never increase by 2 or decrease to a value other than 1.
In this paper, we pose the following research question: Can we develop a Whittle index-based online algorithm for RMABs with non-stationary transition kernels?
@@
  • Algorithm Design. The challenge for designing online learning algorithms for RMABs is to incorporate the computationally efficient class of policies such as Whittle index policy into an adaptive process. We design a sliding window-based online Whittle index policy for non-stationary RMABs (see Algorithm 1). We model non-stationarity of transition kernels of arm $n$ by using a total variation budget $V_{n}$ which is an upper bound of the sum of the total variational distance $\tilde{V}_{n}$. To estimate the budget $V_{n}$, we utilize a Bandit-over-Bandit approach [33, 21], in which $V_{n}$ is selected from a finite set of possible values. Based on the estimated $V_{n}$, the Whittle index is predicted by using a sliding window and upper confidence bound approaches. Moreover, our algorithm takes into account the sparsity of the transition kernels. This significantly simplifies the complexity of optimization and helps to predict the transition kernels accurately.
  • Dynamic Regret Analysis. We rigorously characterize an upper bound on the dynamic regret of our algorithm. Our paper is the first to provide dynamic regret for the online learning of Whittle index under non-stationary environments. It is difficult to analyze dynamic regret of an online policy under non-stationary environments. It is even more difficult for RMABs. [30] overcame this challenge by analyzing the regret for stationary environment using the Lagrangian relaxed form of the problem and its solution. In this paper, we extend the regret analysis to (i) non-stationary environments and (ii) to a stronger version of regret by directly analyzing the main problem and its solution, instead of the Lagrangian form. Our policy can achieve dynamic regret of $\tilde{O}(T^{2/3}\tilde{V}^{1/3}+T^{4/5})$ for large system size when RMAB is indexable and has a global attractor point (see Theorem 3 & Remark 1).
  • Simulation Results. Our simulation results (see Table 1 & Fig. 1) show that our algorithm achieves lower regret in practice compared with the UCWhittle policy [30], WIQL policy [34], and a uniformly randomized policy [3] baselines.
@@ 2 Related Work
Offline Whittle Index Policy for RMABs: Whittle’s seminal work [1] introduced a heuristic policy for the infinite-horizon RMAB problem, known as the Whittle index policy. Motivated by Whittle’s work, many subsequent works have applied the Whittle index framework to different resource allocation problems [28, 4, 5, 29, 12, 3, 6, 35] by modeling them as RMABs.
Online Learning of Whittle Index: Multiple works [36, 37, 34] have proposed Q-learning algorithms to compute Whittle Index. Authors in [38] proposed NeurWIN and [39] proposed DeepTOP to compute Whittle index by using neural networks. These prior works did not provide any regret guarantees for their policy. In [40], the authors develop an online Whittle algorithm with static regret guarantees compared to the best fixed Whittle index policy. [30] is the first to provide the regret analysis for the online learning of Whittle index with unknown transition kernels. However, [30] consider a stationary environment. In [30], authors analyzed regret of UCWhittle by using Lagrangian relaxed form of the RMAB problem. We, in this paper, propose an online learning of Whittle index for non-stationary transition dynamics, with provable regret bounds. To the best of our knowledge, this is the first work to provide dynamic regret analysis of an online Whittle index-based policy for RMABs with non-stationary transitions.
@@ 3 Problem Setting
We consider an episodic RMAB problem with $N$ arms and an unknown non-stationary environment. Each arm $n\in[N]$ is associated with a unichain MDP denoted by a tuple $(\mathcal{S},\mathcal{A},P_{n,t},r_{n})$ at every episode $t$, where the state space $\mathcal{S}$ is finite, $\mathcal{A}=\{0,1\}$ is a set of binary actions, $P_{n,t}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ is the transition kernel of arm $n$ with $P_{n,t}(s^{\prime}|s,a)$ being the probability of transitioning to state $s^{\prime}$ from state $s$ by taking action $a$ in episode $t$, and $r_{n}(s,a)$ is the reward function for arm $n$ when the current state is $s$ and the action $a$ is taken. The total number of episodes is $T$ and each episode itself consists of $H$ time slots. We consider that the transition kernels $P_{n,t}$ are unknown and non-stationary, i.e., $P_{n,t}$ can change across episodes $t\in[T]$.
A decision maker (DM) determines what action to apply to each arm at a decision time $h\in[H]$ of an episode $t\in[T]$ under the instantaneous activation constraint that at most $M$ arms can be activated. The action taken by the DM in episode $t$ is described by a deterministic policy $\pi_{t}:\mathcal{S}^{N}\mapsto\mathcal{A}^{N}$ which maps a given state $(s_{1},s_{2},\ldots,s_{N})\in\mathcal{S}^{N}$ to an action $(a_{1},a_{2},\ldots,a_{N})\in\mathcal{A}^{N}$. The corresponding expected discounted sum of rewards in episode $t$ is given by
@@
where $\Pi$ is the set of all causal policy $\pi_{t}:\mathcal{S}^{N}\mapsto\{0,1\}^{N}$.
@@ 3.1 Lagrangian Relaxation
Because the main problem described in equation 2 is intractable, we relax the per-time slot constraint and use the Lagrangian defined below:
@@
The Lagrangian problem described in equation 3 enables us to decompose the combinatorial decision problem equation 2 into a set of $N$ independent Markov decision process for each arm:
$U\left(\pi_{n,t},P_{n,t},\lambda\right)=\max_{\pi_{n,t}\in\Pi_{n}}\mathbb{E}\left[\sum_{h=1}^{H}\gamma^{h-1}\bigg(r_{n}(s_{n,h,t},a_{n,h,t})-\lambda a_{n,h,t}\bigg)\bigg|\pi_{n,t},P_{n,t}\right],$ (4)
where $\pi_{n,t}^{*}$ is the optimal solution that maximizes equation 4 from the set of all causal policies $\Pi_{n}$
@@ 3.2 Whittle Index Policy
Given $\lambda$, we denote by $\phi_{n}(\lambda)$ the set of states for which it is optimal not to activate the arm. The set $\phi_{n}(\lambda)$ is given by $\phi_{n}(\lambda):=\{s\in\mathcal{S}:Q_{n,\lambda,t}(s,0)>Q_{n,\lambda,t}(s,1)\}$, where the action value function $Q_{n,\lambda,t}(s,a)$ associated with Bellman optimality equation for equation 4 is
@@
and the value function $V_{n,\lambda,t}(s)$ associated with Bellman optimality equation for equation 4 is
@@
Definition 1 (Indexability)
An arm is said to be indexable if the set $\phi_{n}(\lambda)$ increases monotonically as $\lambda$ increases from 0 to $\infty$. A restless bandit problem is said to be indexable if all arms are indexable.
Definition 2 (Whittle Index)
@@
Whittle Index Policy activates at most $M$ arms out of $N$ arms with highest Whittle indices. However, as we can observe from equation 7, we can compute Whittle index if we know the transition kernel $P_{n,t}$ of every episode $t\in[T]$. Next, we model how transition kernels change over every episode.
3.3 The Transition Kernel Model
Non-Stationarity: In this section, we model the transition kernels for our non-stationary RMAB setting. We assume that the transition kernels $P_{n,t}$ may drift at varying rates across different arms $n\in[N]$ with the constraint that the total variation distance between transition kernels of two consecutive episodes is bounded from above by
$\max_{(s,a)\in\mathcal{S}\times A}\sum_{s^{\prime}\in\mathcal{S}}\bigg|P_{n,t}(s^{\prime}|s,a)-P_{n,t-1}(s^{\prime}|s,a)\bigg|\leq\frac{V_{n}}{T},$ (8)
where $V_{n}$ is the total variation budget across the entire $T$ episodes. The total variation budget $V_{n}$ represents the total non-stationarity in arm $n$ across the entire horizon, and is a standard quantity used for analzing dynamic regret in online learning literature [18, 19].
Sparsity: In many applications, the probability transition kernels are sparse - meaning that many state transitions are not possible under certain actions. To model this we introduce $\mathcal{S}_{0}(s,a)$ as the set of all states $s^{\prime}\in\mathcal{S}$ such that the probability to transit from state $s\in\mathcal{S}$ to state $s^{\prime}\in\mathcal{S}$ given action $a\in\mathcal{A}$ is always 0, i.e.,
$\mathcal{S}_{0}(s,a)=\{s^{\prime}\in\mathcal{S}:P_{n,t}(s^{\prime}|s,a)=0,\forall t\}.$
The sets $\mathcal{S}_{0}(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$ represents the sparsity of transition kernels for arm $n$. Our proposed algorithm can utilize this sparsity to provide better regret bounds. However, even in the absence of any sparsity, we are able to guarantee sublinear dynamic regret.
The DM is assumed to know the parameter $\mathcal{S}_{0}(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. In next section, we develop our Algorithm 1 that (i) learns the total variation budget and the probability transition kernels, and (ii) uses them to compute the Whittle Index to pick approximately optimal policies in each episode. In the next section, we discuss how we obtain our online algorithm.